# A Parameter Update Balancing Algorithm for Multi-task Ranking Models in Recommendation Systems

### v1

@@ Abstract
Multi-task ranking models have become essential for modern real-world recommendation systems. While most recommendation researches focus on designing sophisticated models for specific scenarios, achieving performance improvement for multi-task ranking models across various scenarios still remains a significant challenge. Training all tasks naïvely can result in inconsistent learning, highlighting the need for the development of multi-task optimization (MTO) methods to tackle this challenge. Conventional methods assume that the optimal joint gradient on shared parameters leads to optimal parameter updates. However, the actual update on model parameters may deviates significantly from gradients when using momentum based optimizers such as Adam, and we design and execute statistical experiments to support the observation. In this paper, we propose a novel Parameter Update Balancing algorithm for multi-task optimization, denoted as PUB. In contrast to traditional MTO method which are based on gradient level tasks fusion or loss level tasks fusion, PUB is the first work to optimize multiple tasks through parameter update balancing. Comprehensive experiments on benchmark multi-task ranking datasets demonstrate that PUB consistently improves several multi-task backbones and achieves state-of-the-art performance. Additionally, experiments on benchmark computer vision datasets show the great potential of PUB in various multi-task learning scenarios. Furthermore, we deployed our method for an industrial evaluation on the real-world commercial platform, HUAWEI AppGallery, where PUB significantly enhances the online multi-task ranking model, efficiently managing the primary traffic of a crucial channel.
@@ I Introduction
Multi-task learning (MTL) is a powerful technique widely applied to various recommendation systems [1, 2, 3, 4, 5, 6], including e-commerce, social media, news/video feeds, and online advertising. For instance, in e-commerce, a ranking model can be trained on multiple related tasks, such as predicting user ratings, click-through rate, and purchase likelihood. This model learns shared representations of users and items, which can be used to make personalized recommendations. Similarly, in social media, an MTL model can be trained on tasks such as predicting user engagement with posts, comments, and likes. MTL has demonstrated significant potential for enhancing the accuracy of recommendation systems and user experience.
Current multi-task models often share some parameters across all tasks to learn shared representations and then use task-specific parameters to predict multiple labels. However, the intricate and competing correlations among tasks may result in inequitable learning if all tasks are naively trained together. This problem is also known as the seesaw problem [7], where improving performance in one task comes at the expense of other tasks. A major reason behind this issue is conflicting gradients [8], meaning the gradients of different task are not well aligned. As a result, following the average gradient direction can be detrimental to the performance of specific tasks. To address this issue, many multi-task optimization methods have been conducted to alleviate this problem [9, 10, 11, 12, 13, 14]. As shown in Figure LABEL:motivation(a), these methods try to combine per-task gradients (black arrow) into an optimal joint direction (blue arrow). Then, they employ an optimizer to compute parameter updates $\Delta\boldsymbol{\theta}$ (red arrow) to modify model parameters. Although existing MTO methods could alleviate seesaw problem to some extend, they are all based on gradient level tasks fusion or loss level tasks fusion, and weighted loss will converge to gradient balancing with regard to shared parameter updating. It is worth noting that the actual updates of model parameters may deviate significantly from their gradients [6], especially when employing momentum based optimizers like Adam and AdaGrad, etc.
In this paper, we propose a Parameter Updating Balancing method, denoted as PUB, to overcome the seesaw problem in optimizing multi-task ranking model with shared parameters. As illustrated in Figure LABEL:motivation(b), unlike traditional methods that balance task gradients, PUB aims to effectively identify the optimal joint parameter update based on task updates. In this way, PUB is able to overcome the limitations of conventional methods mentioned above. As shown in Figure LABEL:fig:toy, we conducted a toy experiment with Adam optimizer to intuitively understand the differences between PUB and several state-of-the-art (SOTA) methods. Our contributions are summarized as follows:
@@ III-A Root Cause and Evidence
We employed the Chi-square test of independence to determine if there is a significant association between $x$ and $y$. The test revealed a statistically significant association between the two variables, $\chi^{2}(1)=13.67,p<0.001$. The 70% of the results demonstrate a reverse correlation between average AUC and Diff, i.e. lower Diff has high probability of causing higher overall multi-task performance.
@@ IV Preliminary
$T$ Total training step number
$t$ Current training step
@@ V Method
Based on Section III, we know that without gradients of other tasks, update obtained by task gradient ($i.e.\Delta\boldsymbol{\theta}_{i}=f(\boldsymbol{g}_{i})$) is sufficient for the multi-task ranking. In this section, we detailed describe PUB an efficient method to obtain optimal update. In this section, we first describe PUB algorithm in general to give a overall picture. Then we formalize the update balancing problem as a convex optimization problem. Lastly, we describe the method to efficient approximate the convex optimization problem.
@@ V-A PUB Algorithm
When updating the momentum, we simply use the same weights as aggregating gradient. On the one hand, this trick can reduce complexity and save training time; on the other hand, it does not appear to have a negative impact in our experiments. We leave a theoretical analysis of this trick in the future work. In our practice, PUB updates moment vectors after updating parameters, but people can seem it as a hyperparameter and tune it according to actual situation.
@@
Theoretically, PUB can be integrated with different optimizers and UMMs easily. Because above analysis does not make any assumption or constrain on the method to get $\Delta\theta_{i}$. PUB has two differences when integrating different optimizers, computing task updates $\Delta\theta_{i}$ and updating moment vectors. To demonstrate the flexibility of PUB in a low cost way, we have conducted experiments on PUB with Adam and some UMMs and left the extension research for future work.
@@ VI-B Two tasks experiments in Recommendation Ranking
In real-world applications, there may be some specific requirements besides alleviating the seesaw problem, such as improving CTCVR more for online advertising platforms. UMMs are commonly used methods to satisfy such special requirements. Fortunately, PUB is flexible and can integrate UMMs easily to accommodate such real-word special requirements. We will assess its capability in the upcoming experiment. It should be noted that conventional MTO methods are unable to integrate UMMs, since they are all based on gradient level tasks fusion or loss level tasks fusion.
@@
PUB with UMMs. To demonstrate flexibility of our proposed method, we evaluate PUB with two SOTA update manipulation methods, Clippy [6] and AdaTask [28]. We re-implement Clippy using Adam according to its paper based on the source code of Fairseq [40]. It should be noted that current UMMs are based on certain assumptions, so they may be ineffective in some datasets which do not meet the assumptions. Besides, these methods often require significant effort to adjust hyperparameters and certainly result in larger training cost. Thus, to save costs, we only comprehensively evaluate PUB in the AliExpress_ES dataset. We search tasks weights every 10 steps in training. Table IV demonstrates the experimental results of PUB with different UMMs. We did not conduct experiments that integrating UMMs with conventional MTO methods, since these methods are hard to combine UMMs because of their fundamental design.

### v2

@@ Abstract
Multi-task ranking models have become essential for modern real-world recommendation systems. While most recommendation researches focus on designing sophisticated models for specific scenarios, achieving performance improvement for multi-task ranking models across various scenarios still remains a significant challenge. Training all tasks naïvely can result in inconsistent learning, highlighting the need for the development of multi-task optimization (MTO) methods to tackle this challenge. Conventional methods assume that the optimal joint gradient on shared parameters leads to optimal parameter updates. However, the actual update on model parameters may deviates significantly from gradients when using momentum based optimizers such as Adam, and we design and execute statistical experiments to support the observation. In this paper, we propose a novel Parameter Update Balancing algorithm for multi-task optimization, denoted as PUB. In contrast to traditional MTO method which are based on gradient level tasks fusion or loss level tasks fusion, PUB is the first work to optimize multiple tasks through parameter update balancing. Comprehensive experiments on benchmark multi-task ranking datasets demonstrate that PUB consistently improves several multi-task backbones and achieves state-of-the-art performance. Additionally, experiments on benchmark computer vision datasets show the great potential of PUB in various multi-task learning scenarios. Furthermore, we deployed our method for an industrial evaluation on the real-world commercial platform, HUAWEI AppGallery, where PUB significantly enhances the online multi-task ranking model, efficiently managing the primary traffic of a crucial channel. Our code is released at https://github.com/yjdy/Multitask-Optimization-Recommendation-Library.
@@ I Introduction
Multi-task learning (MTL) is a powerful technique widely applied to various recommendation systems [1, 2, 3, 4, 5, 6], including e-commerce, social media, news/video feeds, and online advertising. For instance, in e-commerce, a ranking model can be trained on multiple related tasks, such as predicting user ratings, click-through rate, and purchase likelihood. Similarly, in social media, an MTL model can be trained on tasks such as predicting user engagement with posts, comments, and likes.
MTL has demonstrated significant potential for enhancing the accuracy of recommendation systems and user experience. Current multi-task models often share some parameters across all tasks to learn shared representations and then use task-specific parameters to predict multiple labels. However, the intricate and competing correlations among tasks may result in inequitable learning if all tasks are naively trained together. This problem is also known as the seesaw problem [7], where improving performance in one task comes at the expense of other tasks. A major reason behind this issue is conflicting gradients [8], meaning the gradients of different task are not well aligned. As a result, following the average gradient direction can be detrimental to the performance of specific tasks. To address this issue, many multi-task optimization methods have been conducted to alleviate this problem [9, 10, 11, 12, 13, 14]. As shown in Figure LABEL:motivation(a), these methods try to combine per-task gradients (black arrow) into an optimal joint direction (blue arrow). Then, they employ an optimizer to compute parameter updates $\Delta\boldsymbol{\theta}$ (red arrow) to modify model parameters. Although existing MTO methods could alleviate seesaw problem to some extend, they are all based on gradient level tasks fusion or loss level tasks fusion, and weighted loss will converge to gradient balancing with regard to shared parameter updating. It is worth noting that the actual updates of model parameters may deviate significantly from their gradients [6], especially when employing momentum based optimizers like Adam and AdaGrad, etc.
In this paper, we propose a Parameter Updating Balancing method, denoted as PUB, to overcome the seesaw problem in optimizing multi-task ranking model with shared parameters. As illustrated in Figure LABEL:motivation(b), unlike traditional methods that balance task gradients, PUB aims to identify the optimal joint parameter update based on task updates. In this way, PUB is able to overcome the limitations of conventional methods mentioned above. As shown in Figure LABEL:fig:toy, we conducted a toy experiment with Adam [15] to intuitively understand the differences between PUB and several state-of-the-art (SOTA) methods. Our contributions are summarized as follows:
@@ III-A Root Cause and Evidence
We employed the Chi-square test of independence to determine if there is a significant association between $x$ and $y$. The test revealed a statistically significant association between the two variables, $\chi^{2}(1)=16.99,p<0.001$. The 67.36% of the results demonstrate a reverse correlation between average AUC and Diff, i.e. lower Diff has high probability of causing higher overall multi-task performance.
@@ IV Preliminary
$T$, $t$ Total and current training step
@@ V Method
Based on Section III, we know that without gradients of other tasks, update obtained by task gradient ($i.e.\Delta\boldsymbol{\theta}_{i}=f(\boldsymbol{g}_{i})$) is sufficient for updating parameters. This section details PUB, an efficient method for achieving optimal updates. We first provide a general overview of the PUB algorithm, followed by a formalization of the update balancing problem as a convex optimization issue. Finally, we outline a method for efficiently approximating this convex optimization problem.
@@ V-A PUB Algorithm
When updating the momentum, we simply use the same weights as aggregating updates. On the one hand, this trick can reduce complexity and save training time; on the other hand, it does not appear to have a negative impact in our experiments. We leave a theoretical analysis of this trick in the future work. In our practice, PUB updates moment vectors after updating parameters, but people can seem it as a hyperparameter and tune it according to actual situation.
@@
Theoretically, PUB can be integrated with different optimizers and UMMs easily. Because above analysis does not make any assumption or constrain on the method to get $\Delta\theta_{i}$. PUB only has two differences when integrating different optimizers, computing task updates $\Delta\theta_{i}$ and updating moment vectors. To demonstrate the flexibility of PUB in a low cost way, we have conducted experiments on PUB with Adam and some UMMs and left the extension research for future work.
@@ VI-B Two tasks experiments in Recommendation Ranking
PUB with UMMs. In real-world applications, there may be some specific requirements besides alleviating the seesaw problem, such as improving CTCVR more for online advertising platforms. UMMs are commonly used methods to satisfy such special requirements. Conventional MTO methods are unable to integrate UMMs, since they are all based on gradient level tasks fusion or loss level tasks fusion. Fortunately, PUB is flexible and can integrate UMMs easily to accommodate such real-word special requirements. We will assess its capability in the upcoming experiment.
@@
To demonstrate flexibility of our proposed method, we evaluate PUB with two SOTA update manipulation methods, Clippy [6] and AdaTask [29]. We re-implement Clippy using Adam according to its paper based on the source code of Fairseq [40]. It should be noted that current UMMs are based on certain assumptions, so they may be ineffective in some datasets which do not meet the assumptions. Besides, these methods often require significant effort to adjust hyperparameters and certainly result in larger training cost. Thus, to save costs, we only comprehensively evaluate PUB in the AliExpress_ES dataset. We search tasks weights every 10 steps in training. Table IV demonstrates the experimental results of PUB with different UMMs. We did not conduct experiments that integrating UMMs with conventional MTO methods, since these methods are hard to combine UMMs because of their fundamental design.